OpenStudy (anonymous):

How do you find the slope of a line if you have 2 points on the line?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

subtract the points and put y/x

OpenStudy (amistre64):

(3,6) -(1,7) ------ (2,-1); slope = y/x = -1/2

OpenStudy (amistre64):

the text i believe says: the change in y ------------- the change in x

OpenStudy (amistre64):

\[\frac{y-y_o}{x-x_o}\]

OpenStudy (amistre64):

i just keep better track of then when I stack and subtract